### v2

# Slow is Fast! Dissecting Ethereum’s Slow Liquidity Drain Scams

Abstract
We identify the slow liquidity drain (SLID) scam, an insidious and highly profitable threat to decentralized finance (DeFi), posing a large-scale, persistent, and growing risk to the ecosystem. Unlike traditional scams such as rug pulls or honeypots (USENIX Sec’19, USENIX Sec’23), SLID gradually siphons funds from liquidity pools over extended periods, making detection significantly more challenging. In this paper, we conducted the first large-scale empirical analysis of 319,166 liquidity pools across six major decentralized exchanges (DEXs) since 2018. We identified 3,117 SLID affected liquidity pools, resulting in cumulative losses of more than US\$103 million. We propose a rule-based heuristic and an enhanced machine learning model for early detection. Our machine learning model achieves a detection speed 4.77 times faster than the heuristic while maintaining 95% accuracy. Our study establishes a foundation for protecting DeFi investors at an early stage and promoting transparency in the DeFi ecosystem.
1 Introduction
Decentralized Finance (DeFi) [1, 2, 3] has emerged as an innovative landscape with the potential to revolutionize the global financial system [4, 5]. DeFi applications aim to replace traditional financial services by eliminating intermediaries and promoting transparency and permissionlessness through blockchain technology [6].
A pivotal moment for DeFi occurred with the rise of Uniswap [7], the leading decentralized exchange (DEX) on Ethereum. Uniswap introduced the Automated Market Maker (AMM) [8, 9, 10], enabling users to create liquidity pools (LPs) [^1] We use the abbreviation LP to denote liquidity pool, contrasting with sources [1, 11] that define it as liquidity provider. using their newly created tokens and facilitating direct asset exchanges without reliance on third parties like centralized exchanges (CEXs). Subsequent DEXs, such as Curve [12] and SushiSwap [13], further attracted investors by offering a broad range of cryptocurrencies, including SHIB [14] and PEPE [15]. The market value of DEXs [16, 17, 18] surged from \$587K in 2018 to \$7.71B during the “DeFi summer” of 2020 and eventually surpassed \$100B in 2021 [19]. To date, DEXs handle billions of dollars in daily trading volume [20].
Decentralized exchanges have unsurprisingly become prime targets for financial scams [21, 22]. De.fi reported [23] a staggering \$1.39 billion loss in Ethereum’s DeFi sector due to scams. These scams often involve the rapid creation of liquidity pools, allowing fraudsters to retain full control over their value. Once users invest, they find themselves unable to sell or recover their funds. Two infamous examples include the rug pull [24], where developers abruptly abandon a project after withdrawing liquidity, and the honeypot [25], which lures users by permitting deposits but blocking withdrawals.
Although advancements in scam detection and increased user awareness have reduced the prevalence of such straightforward scams, fraudsters are now resorting to more sophisticated schemes designed to bypass detection mechanisms and exploit unsuspecting users.
Consider the case of a GoPlus report that flagged the Emily token as a potential rug pull due to the presence of an “unlimited mint function” [26]. However, upon closer examination, the token did not display the typical characteristics of a rug pull. This observation demanded further scrutiny, leading to the discovery of a far more insidious and pervasive threat within DeFi – a new and covert scam which we term the slow liquidity drain (SLID) scam. This phenomenon represents a sophisticated evolution in fraudulent tactics, exploiting the trust and complexity of decentralized finance to remain hidden. Unlike traditional DEX scams that exhibit clear characteristics (e.g., abrupt withdrawal, sale restriction), the SLID scam operates without disrupting the normal functionality of the liquidity pool. Instead, it gradually siphons assets through a slow but calculated profit-taking strategy. This stealthy approach enables the scam to blend with legitimate pools, making detection more challenging. As a result, investors may experience prolonged losses without immediate realization.
Our analysis exposes the substantial fraudulent profits generated by SLID scams. Between 2018 and 2022, SLID scams siphoned off over \$64.33M in realized illicit gains, accounting for 14.3% of the total decentralized exchange (DEX) scam losses on Ethereum and Binance Chain combined (estimated at \$400M [27]). This striking figure underscores the broad impact of SLID scams within the DeFi ecosystem. Extending the scope to cover 2018 through 2024, these scams have inflicted a cumulative loss of \$103M, including realized and unrealized profits, marking them as a persistent threat to DeFi. Furthermore, our analysis of the dataset collected from Ethereum [28] reveals a yearly increase in the number of SLID scams since their emergence in 2018 (Figure 9). This trend highlights the urgency of raising awareness about SLID scams to mitigate the growing threat they pose to DeFi ecosystems.
We progress by addressing the following questions.
  • $\bullet$
    RQ1. What are the common features of SLID scams? Currently, SLID scams lack formal recognition. Their defining characteristics remain unexplored. We aim to establish a clear definition of these scams, aiding in the identification of similar patterns in SLID liquidity pools and facilitating their distinction from legitimate ones.
  • $\bullet$
    RQ2. Can we build and evaluate a rule-based detector to identify SLID scams? Using the identified characteristics of SLID scams, we aim to design a rule-based heuristic to detect SLID liquidity pools on Ethereum. This heuristic will be evaluated in conjunction with external data from legitimate liquidity pools. Additionally, we aim to gain deeper insights into SLID scams by analyzing cases with unexpected outcomes (i.e., false positives, false negatives).
  • $\bullet$
    RQ3. Can we use the heuristic to automatically detect SLID scams at a large scale? Based on RQ2, we aim to apply the heuristic at scale to identify and label all SLID scams across the six largest Ethereum DEXs. However, the heuristic’s dependency on a long historical range of transaction data introduces latency in flagging SLID scams. To address this limitation, we also aim to develop a more proactive method capable of providing early and rapid warnings.
We first established our initial ground truth dataset by identifying community-reported rug pull cases that later failed to trigger the proposed rug pull detectors and exhibited characteristics similar to our research motivation. This process resulted in the identification of 71 liquidity pools that matched these descriptions. We then conducted a pilot study to analyze common scam behaviors shared by this initial ground truth dataset (§5). Based on the insights extracted from this analysis, we designed an efficient rule-based heuristic capable of identifying SLID liquidity pools by examining their DEX activity records and smart contract security features. This heuristic was tested using legitimate liquidity pools, with the evaluation outcomes discussed in detail (§6).
Figure 1: The workflow of our SLID scam: ① create a liquidity pool; ② add tokens (base & paired); ③ avoid burning initial shares/tokens; ④ marketing campaign; ⑤ invest funds and buy tokens; ⑥ normal activities (buy orders); ⑦ slowly withdraw/sell funds; ⑧ liquidity drop; ⑨ money loss.
Next, we applied the proposed rule-based heuristic to 319,166 Ethereum liquidity pools across the six largest DEXs on Ethereum (§7). This resulted in the flagging of 3,117 SLID scams, which collectively generated a profit of \$103.2M on these DEXs as of July 13, 2024. This analysis introduced the first-ever ground truth dataset of confirmed SLID scams. We then analyzed the flagged instances through three key dimensions: vitality, trend analysis, and profit, providing deeper insights into the tactics and the overall impact of SLID scams. Leveraging this expanded ground truth, we applied three ensemble machine learning (ML) algorithms (§7.3) of Random Forest, XGBoost, and Logistic Regression for early detection of SLID scams, capable of identifying potential scams with a shorter required DEX activity history that can detect SLID scam 4.77 times faster than the heuristic while maintaining high detection accuracy of 95%.
In summary, we made the following key contributions:
  • $\bullet$
    We presented the first formal definition of the SLID scam in DEX environments. We conducted an efficient pilot study to explore its characteristics and design a rule-based heuristic. We then evaluate the heuristic and discuss insights behind the different true and false positive detections.
  • $\bullet$
    We applied our heuristic to six major DEXs on the Ethereum network, identifying 3,117 liquidity pools exhibiting SLID behaviors (with responsible disclosure). We produced the first ground truth dataset for SLID scams, publicly released at https://anonymous.4open.science/r/SLID-sourcecode-0DA3, and enabled analysis across vitality, profit, and trend metrics.
  • $\bullet$
    To enable early warning, we further applied ML algorithms using our ground truth dataset. We successfully reduced detection time by 4.77x while maintaining 95% accuracy.
[…]
2 Technical Warm-Ups
Ethereum tokens. Ethereum [28, 29] is the second leading blockchain platform after Bitcoin [30]. Ethereum introduced smart contracts and enabled the automatic execution of code. Users can create tradeable digital assets on-chain, known as tokens. These tokens serve multiple purposes, including fueling decentralized applications (DApps), acting as shares/stakes in projects, or representing ownership in organizations. Most tokens on Ethereum adhere to ERC20 [31], which defines a set of functions and interfaces. The standard ensures interoperability, allowing smooth token transfers among users.
Ethereum accounts are tied to unique key pairs and are categorized into two types [32]: (i) Externally owned accounts (EOAs) store tokens and facilitate interactions with other accounts; (ii) contract accounts are associated with smart contracts to provide specialized on-chain services.
DEXs and liquidity pool.  Uniswap [33], the unicorn DEX on Ethereum, is the vital DApp that has significantly contributed to the growth of the DeFi ecosystem. Their AMM infrastructure [9] enables users to list their newly created tokens for trading by deploying a liquidity pool. A liquidity pool is a smart contract on the DEX that holds two or more tokens, typically including at least one well-known token with independent value (such as ETH, USDT, or USDC), referred to as the base token. The other token is another base token or a newly introduced token, known as the paired token.
When initializing a liquidity pool, the owner must deposit a specific amount of both tokens. The value of this LP is determined by the total USD value of the base token it holds, while the price of the paired token is positively correlated to the pool’s overall value, following the provided AMM equations [34]. Trading the base token for the paired token is termed a buy, while trading in the opposite direction is a sell. These swap actions impact the pool’s dynamics: a buy increases the price of the traded token as more base tokens flow in, whereas a sell reduces the price as tokens are withdrawn.
Any user can contribute to the pool’s liquidity via deposit with a certain amount of both tokens to earn incentives later or through withdraw to remove their previously provided liquidity. All activities are publicly recorded on-chain and are referred to as DEX orders or DEX activities of that pool.
Legitimate LPs conduct the action of burning (i.e., permanent locking) initial LP tokens after deployment. This process removes the pool owner’s ability to control the liquidity, preventing them from taking unethical actions with investor funds, such as draining the pool. Burning LP tokens is widely regarded as an industry standard for establishing trustworthiness and transparency in DeFi space [33, 1, 2, 35].
Rug pull.  The most common DEX scam leveraging liquidity pools is rug pull [24]. The scammer creates (so that owns) the scam pool and extracts all the base tokens. There are two basic operations through which scammers extract profits:
  • $\bullet$
    Sell: first, sell a large inflated amount of paired tokens (they owned) in exchange for all pool’s base tokens;
  • $\bullet$
    Withdraw: withdraw the liquidity they initially provided, “pulling” out all of the base tokens from the pool.
When scammers engage in sell or withdraw activities and the value of the assets they retrieve exceeds their initial investment in the scam pool, the excess amount is defined as the scammer’s realized profits. Correspondingly, unrealized profits represents the remaining value of assets within the scam pool that still belong to the scammer but have not yet been converted into realized profits via sell or withdraw actions.
Those sudden profit-taking actions caused a massive drop in the paired token price and the pool’s trading volume, serving as common signals for detecting rug pull.
Honeypot. Another common DEX scam is Honeypot [25]. The scammer maliciously designs the token’s smart contract to allow users to easily purchase the token but block them from selling or withdrawing it. The contract may include restrictions such as preventing sell transactions, allowing withdrawals only for whitelisted addresses, or imposing exorbitant transaction fees that make it impossible to retrieve funds. The scammer then exploits the liquidity pool through sell and withdraw actions, similar to the mechanics of a rug pull scam.
We defer more related work in Appendix A.
All the stated fraudulent behaviors share common characteristics regarding the liquidity pool: (i) the liquidity pool tokens (LPTs), which represent the initial liquidity [36], are not burned, and (ii) the ownership of the paired token remains entirely monopolized by the pool creator. Without implementing an ethical safeguard (i.e., burning initial LPTs), liquidity providers, including those involved in rug pulls, honeypots, and SLID scams (as identified in our findings), confront no financial risk (proven in Appendix B) for engaging in fraudulent activities. This enables them to directly drain assets from users’ investments, perpetuating the cycle of exploitation.
3 Motivation Example
This section discusses the behavior of Token $\hat{A}$ (all the collected data, including token names, addresses, and liquidity pool details, are anonymized for privacy & ethical reasons) with token address “0x4e84…a17c8”, created in August 2021, which serves as a motivational example for our study. This token was launched with a massive marketing campaign highlighting its high potential for generating profits for investors.
We have identified the liquidity pool of $\hat{A}$ as a SLID scam due to the deceptive actions of its owners. They slowly exploited the liquidity pool through a mixed sequence of token buying and selling on Uniswap while maintaining a stable price and trading volume during the initial months after its launch. The gradual profit-taking persisted for an extended period before the token owner executed a significant rug pull a year later. We briefly describe the sequence of actions taken by the owners in this cryptocurrency’s SLID scam.
① Attracting potential victims. To rapidly attract users, $\hat{A}$’s advertising campaign targeted buyers by promising huge token profits if purchased early. The owner engaged the community by promoting the token via their X account (formerly Twitter). They attracted 78.5K followers and collaborated with 20+ well-known crypto influencers. The campaign successfully drew 5,326 investors to purchase $\hat{A}$ on Uniswap within the first 70 days from the token’s LP deployment.
② Maintaining the pool and slowly taking profit.  The token’s liquidity pool was deployed on August 4th, 2021, with an initial owner liquidity deposit of \$19K. However, the liquidity pool for $\hat{A}$ was deployed without burning the owner’s initial tokens. This creates a persistent risk, as the owner of $\hat{A}$ can drain the liquidity pool and seize all funds invested by users to purchase $\hat{A}$ at any time. This is precisely what occurs in rug pull and honeypot scenarios (as described in §2).
We observed $\hat{A}$’s owner DEX activities (cf. Figure 2) over 70 days after deployment, including price and trading volume. During this period, the owner was engaged in a sequence of mixed activities, including buying and selling. Although these actions seem common and legitimate, our evaluation reveals that the selling amount consistently exceeded the buying amount (in paired tokens). Specifically, the total selling volume was nearly ten times larger, with the owner purchasing only $2.38\times 10^{15}$ tokens while selling $3.21\times 10^{16}$ tokens of $\hat{A}$ in total. This indicates that the owner possesses a significant number of tokens and attempts to incorporate this inflated amount into their selling orders. $\hat{A}$’s owner executed 139 buy and 136 sell actions, but with a nearly double volume gap that was worth \$783K and \$1.4M, respectively. This resulted in a huge profit for the pool owner of \$707K — nearly double the initial investment (of the base token amount the owner used to initially deposit and buy the paired token multiple times) — that comes directly from the investor’s money.
In contrast, in this 70-day period, those 5,326 investors committed with 10,374 buys, 3,898 deposits, 7,322 sells, but only 2 withdraw activities, with a total inflow of nearly \$1.35M (in base token) to the pool. The mix of buy and sell orders created the appearance of legitimacy in the token’s transaction history, helping to avoid the token price crash that would occur if they sold all at once (i.e., 1-day rug pull [37]).
③ Repeating the scam. While maintaining the appearance of token legitimacy, $\hat{A}$’s owner continued to sell only a small portion of the tokens compared to the overall LP value. The smallest and largest sell order values account for only 1% and 16% of the pool’s value. By carefully mixing buy and sell orders and controlling the selling amounts, the built-in scam detectors on DEXs are unable to identify this token as a scam. As of July 13, 2024, the scammer still maintains \$91K worth of liquidity in the pool and has remained inactive since their last recorded sell order on July 31, 2022.
Thus, despite the stable behavior of the token, a series of buying and selling actions (with smaller amounts) occurred at a slow pace, gradually adding an inflated amount of tokens to each sell order from the pool owner’s account. The sequence of actions resembles the profit-taking sell actions seen in rug pull, where an inflated amount is sold to drain the base tokens. This scam strategy can easily avoid investor attention while enabling the token to remain active and attract additional investments over an extended period.
4 Threat Model & Security Goal
Entities in DEXs. A typical DEX involves two entities: the pool owner (short for the owner), represented by the address that deploys the liquidity pool, and the investors/users, who participate in the pool’s activities.
Threat model. An SLID adversary (i.e., the pool owner) may exploit liquidity pools containing user/investor funds in DEXs by employing a gradual profit-taking strategy to: (i) evade the current fraud detection systems; and (ii) maintain user confidence, enabling larger profit over time. An adversary is supposed to have permissions for the following behaviors:
  • $\bullet$
    The pool owner can unethically monopolize the paired token by retaining their initial tokens instead of burning them when creating the liquidity pool. The owner can then extract profits directly from users’ investments/funds (in base tokens) and use profits to buy additional paired tokens.
  • $\bullet$
    The pool owner can perform activities including selling/withdrawing and buying/depositing at a relatively high frequency, each time involving a small amount of tokens. The price and volume for both tokens will not exhibit obvious fraudulent patterns. The owner can extract small portions of profit through these paired activities.
  • $\bullet$
    The pool owner commits to a long-term engagement in such activities, gradually draining funds while maintaining the pool’s appearance to avoid raising suspicion.
Security goals. We have the following three goals:
  • $\bullet$
    identify common features observed in SLID scams by analyzing sufficient real-world public data collected on Ethereum, addressing RQ1;
  • $\bullet$
    design a heuristic capable of detecting SLID fraudulent behavior from the liquidity pool’s visible activities (e.g., owner activity, profit-taking) and capture the occurred SLID fraud among Ethereum DEXs, addressing RQ2;
  • $\bullet$
    enhance detection robustness by enabling real-time identification of SLID scams. Building on the heuristic results, apply ML models to detect SLID scams during the early stages of liquidity pool operations, addressing RQ3.
Figure 2: 70-day scammer activities on Uniswap with token price and trading volume.
5 Common Features of SLID Scam (RQ1)
In this section, we conduct an in-depth investigation into SLID liquidity pools. SLID scam is for the first time identified, where no detection methods can be directly applied. Fortunately, we find that a key feature of SLID is the active engagement of the liquidity pool owner, which creates the illusion of an active and reliable pool while gradually draining profits. Based on this, we generalize (via an empricial study) this single instance into a broader scam pattern that fully captures their distinct features.
5.1 Identification of SLID Pools
Given the lack of prior research on SLID scams, we used rug pulls as an intermediary to collect and analyze data from SLID pools in this pilot study. Our hypothesis was largely based on the similarities between SLID and rug pull scams, particularly in their initialization patterns and profit-taking actions, such as selling and withdrawing liquidity. We also hypothesize that some pools reported as rug pulls may actually represent potential SLID scams.
We first gathered rug pull reports dating back to 2022. Reliable sources for these reports included well-known crypto scam platforms such as Chainabuse [38], De.Fi [39], Chainanalysis [40], and scam warning accounts on X (formerly Twitter), such as ZachXBT [41] and Pocket Universe [42]. Our data collection yields a total of 921 reported rug pulls.
We then designed additional filtering mechanisms to enhance data validity and identify pools displaying SLID characteristics. The reported liquidity pools underwent a filtering process through four conditional layers in the following order:
  • $\bullet$
    Owner profit-checking layer. This layer computes the realized profit of the pool owner based on trading history data and excludes liquidity pools with negative realized profit. Since profit is a vital signal of the scam, this layer ensures the accuracy of the reports by filtering out the falsified reports where the token itself was not fraudulent but reported as a scam. As a result, 23 tokens were excluded from the dataset, specifically those associated with liquidity pools where the owner showed no profit. To be noted, this filtering process may also inadvertently exclude instances of failed scams or pools where the owner benefits through addresses not directly associated with the creator.
  • $\bullet$
    Honeypot layer. As the honeypot is relatively easy to detect during the pool deployment phase, we filtered out instances that are honeypot but reported as rug pull in our collecting process by utilizing GoPlus [43] and De.Fi [39] services. In this layer, we excluded 139 honeypot liquidity pools, which goes against the SLID predetermined characteristic of not having the token’s smart contract interference.
  • $\bullet$
    Mislabelled pools filtering layer. This layer processes the remaining liquidity pool instances and cross-references them with the rug pull detection methods proposed in previous studies [24, 37]. Out of these, 665 instances were confirmed as rug pull and filtered out, while 94 liquidity pools were identified as not being rug pull despite being reported as such on the online resources (claimed in §1).
  • $\bullet$
    Owner action layer. Amongst the 94 remaining instances, we recognize that they may contain various other frauds. To only select scams that align with our motivation example and the study purpose on SLID scams, we selected liquidity pools without the initial LP burned (which pool exists with the risk that the owner has all right to drain the pool at any time after getting investment, discussed at §2) and having at least three DEX activities committed by the owners, including deposit, withdraw, buy, or sell. This filter ensures that the remaining liquidity pool is aligned with our SLID predefined characteristics (§3) of having multiple activities executed by the pool owner. As a result, we excluded 23 liquidity pools and identified 71 pools that passed all four layers as our SLID initial ground truth for RQ1.
5.2 Features Finding of SLID Pools
We then conducted a comprehensive analysis of the filtered 71 liquidity pools and identified three common characteristics (C1, C2, and C3) of SLID pools. For comparison, our analysis included rug pull tokens that had been filtered out by the mislabeled pools filtering layer5.1). We also adopted rug pulls as the benchmark due to the similarities between the two scam types and obtain the distinct features that set SLID scams apart. Our analysis compared LPs from these two scam types across the following three aspects.
Figure 3: 30-day liveliness of rug pull and SLID LPs.
Post-month liveliness analysis.  One of the key metrics that can accurately distinguish SLID scams from rug pulls is post-month liveliness behavior. Cernera et al. [37] found that nearly 60% of the rug pull liquidity pools ended up “inactive” after one day of LP deployment. They further explored LP lifetimes for both scam types over an extended period of one month (cf. Figure 3), providing insights related to SLID scams as well. They revealed that 72.85% of SLID LPs remain active after the first month of deployment, whereas only 10.8% of rug pull LPs persist beyond this threshold.
To further investigate the differences in liveliness between the two scam groups, we compare the average changes in user activity for liquidity pools that remain “active” after the first month. On the first day of the second month, trading activities for the remain-active SLID liquidity pools were stable, showing an average reduction of 38.16% compared to the pool’s previous peak day with the highest recorded activity count. In contrast, the remaining rug pull liquidity pools experienced a dramatic decrease in recorded activities on the first day of the second month, averaging a 95.82% reduction. While other studies on cryptocurrency frauds (e.g., [37] on rug pull, [44] on Ponzi scheme, [45] on NFT rug pull) and detection tools tend to rely on the token’s liveliness as a vital factor for identifying scams, this metric appears to be ineffective in the SLID case due to its strong survival trait.
Our finding in SLID pool’s liveliness led to the following conclusion of SLID characteristic:
  • C1.
    The SLID liquidity pools are more adept at remaining undetected due to their prolonged liveliness. SLID scam pools remain active over a longer duration compared to rug pull despite using similar pool initiation and profit-taking mechanisms, making it nearly impossible to detect based solely on liveliness status.
Scammer activity analysis.  We also analyzed the scammer’s activity patterns in all targeted liquidity pools after the first month to learn the owner’s financial behaviors and motivations behind both fraud types. Given that the scammer is also the owner of these LPs, after victims buy paired tokens and send base tokens in the pool (pool value thus increased), scammers behind both rug pulls and SLID scams typically utilize two mechanisms for profit extraction: $(i)$ withdrawing liquidity from the pool and $(ii)$ selling the inflated token amount they hold externally to drain the liquidity pool.
Rug pull scammers prioritize immediate profit-taking, with 99.7% (663/665) of liquidity pools being entirely drained within the first day, and 100% drained within the first three days. The pool owners, who are also scammers, perform 75.38 DEX activities to their liquidity pool and 15.79 profit-taking actions on average during the first month. Among these profit-taking actions, each pool includes a single, large-amount profit-taking transaction (either selling or withdrawing) that constitutes over 95% of the total pool value. On average, those scammers extract 99.33% of the liquidity pool’s value after the first month, leaving only 0.67% as unrealized profit.
Scammers in SLID pools use contrasting strategies, operating more gradually and performing multiple profit-taking activities with smaller proportions. We observed that none of the SLID pools were completely drained within the first three days after deployment, in stark contrast to the 100% of all pools that were totally drained in rug pull scams during the same period. On average, SLID scammers executed 996.59 total orders during the pool’s first month, which was over 10 times more than the rug pull pool owners. SLID owners had a mean of 423 profit-taking actions, approximately 26 times more frequent than rug pull scammers. Additionally, SLID scammers control the proportion of profit-taking orders in a smaller range. The amount drained value from their profit-taking varies from a minimum of 7.39% to a maximum of 42.93% of the liquidity pool’s total value on average. This controlled proportion enables SLID scammers to evade detection by rug pull detectors, which typically focus on identifying pools with significant exit orders initiated by the owner.
To better demonstrate the differences in scammer actions between two scam types, we further analyzed (Figure 4) the impact of profit-taking activities by scammers in the MIND token [46] (rug pull liquidity pool) and the EMILY token [47] (confirmed inflated mint for rug pull by GoPlus [26] and found in our ground truth of SLID in §5.1) during their first three days. While a single profit-draining action by the MIND token owner completely emptied the liquidity pool and reduced the token’s price to zero, the EMILY token owner conducted multiple small-proportion exits that do not heavily affect the token’s price, thereby retaining future buyers’ confidence. Although both scam types initiate fraud similarly by creating liquidity pools and profiting through draining them via sell or withdraw orders, we observed that variations in execution duration and order proportions lead to distinct strategies for each scam type. From our analysis, we can conclude that:
  • C2.
    Although SLID scams employ a mechanism similar to that of rug pull, they pursue an opposite strategy characterized by a gradual approach, avoiding recognition. Scammers prioritize multiple small-proportion and slow-paced profit extraction activities, making it more difficult to detect, either through observation or by using detectors.
Figure 4: Scammer’s profit-taking activities and its effect on the token price of EMILY (SLID) and MIND (rug pull) liquidity pool in the first three days.
Profit and escalating potential analysis.  Based on the previous observation, the SLID strategy is expected to yield a better profit compared to rug pull by allowing scammers to conceal their activities while gradually attracting more victims. We validate this speculation via a profit-changing analysis between the two scam types. Technically, we focus on the realized profit and the unrealized profit of the LP owner.
We found that the realized profits (the actual profits sold or withdrawn by scammers from the pool) significantly differ between the two scam types. SLID scams demonstrate a superior profit margin: in the first month, while rug pull scammers achieve an average realized profit of 7.5 times their investment through a single, short-term draining action, SLID liquidity pools attain a massively higher profit margin of 10.3 times their investment through multiple small draining activities.
As for unrealized profits, SLID scames also perform a notable gap in profit advantage. Rug pull scammers have nearly zero future profit while completely emptying their liquidity pools, retaining only an average of 0.67% of pool value. In contrast, the SLID pool owner’s drainable assets in the liquidity pool after the first month retain, on average, 156% worth of their initial investment. This allows the SLID scammer, after the first month, to either fully drain the unrealized profit, resulting in an additional 1.56 times profit on top of the realized profit (already taken), or to refrain from withdrawing and let the pool continue to grow.
This observation indicates that by performing a vast number of small-proportion profit-taking orders (§3), SLID scammers manage to attract more victims, resulting in a larger netflow of funds poured into the pool. As a result, although they only take parts of their possible profit in the first month, these scammers already achieve better profit margins than rug pull scammers. Furthermore, a significant portion of assets that remain in the SLID pools retain substantial profit-scaling potential as the pool continues to grow, entrapping more future buyers. Therefore, we conclude that:
  • C3
    SLID scams offer a higher profit margin and sustainability, providing both significant mid-term profits and future profit-scaling potential if undetected. Compared to rug pull, SLID scammers can achieve a substantial growth rate in pool value from their initial investment, making it a more lucrative and enduring scam approach.
[…]
6 Rule-based Detection for SLID Scams (RQ2)
The filters from RQ1 were explicitly designed to separate SLID from rug pull LPs, however, they may perform ineffectively for identifying SLID scams from the population containing various types of liquidity pools. We thus propose a rule-based heuristic designed to effectively detect SLID LPs among the global dataset of Ethereum pools.
6.1 Rule-based Heuristic for SLID Detection
Overview. Our heuristic consists of three validators: Honeypot Validator, profit Validator, and Owner Activity Validator. Each validator targets different characteristics from the extracted data of LPs (cf. RQ1 findings). If one LP is flagged by all three validators (i.e., exhibiting all observed traits), we add it to the initial ground truth dataset as a SLID scam.
Honeypot validator.  This component assesses whether tokens are honeypots, as these are inconsistent with the typical characteristics of an SLID scam. A successful SLID scam often depends on a gradual increase in victim participation and buying volume, fueled by the belief in the token’s legitimacy (see C2). Therefore, it’s crucial for an SLID liquidity pool to display standard initial DEX activities, such as buying and selling, without indications of a honeypot.
Our lightweight Honeypot Validator employs rule-based criteria derived from prior research of [25, 48]. Instead of simply replicating exact detection methods, we apply suggested criteria to specific features extracted from the token’s smart contract and set an upper threshold to identify potential honeypot signals.
In particular, we extracted the following features from the paired token’s smart contract:“buy_tax”, “sell_tax”, “cannot_buy”, “cannot_sell_all”, “slippage_modifiable”, “transfer_pausable”, “personal_slippage_modifiable”, and “trading_cooldown”. These features were evaluated based on specific criteria; each indicates different scenarios in which the pool’s owner can illegally interfere with the user’s activity in the pool for gaining profit. An upper threshold was defined to identify potential honeypots.
Here, we provide an intuitive example: a liquidity pool is flagged as a honeypot if the token’s “buy tax” exceeds 50%, based on tax-related criteria from earlier research [25].
Profit validator.  Profit-related activities reveal the SLID characteristics (§4) from two critical insights (short for I): (i) I-1: the scam likelihood by evaluating realized profit, an essential element of all successful scams; (ii) I-2: the scam likelihood by evaluating the patterns of unrealized profits generated. Thus, we set both realized profit and unrealized profit as our metrics for analyzing the owners’ profit behaviors. Those values can be obtained through the collected DEX orders of liquidity pools (more in §7).
The pool owner’s realized profit can be computed by taking the difference between the total value of their investment in the pool (in USD) and the total value that has been withdrawn over the pool’s lifetime under the base token (detailed equation in Appendix C). We flag a liquidity pool where owners have a positive total realized profit from their activities. This reflects the scam likelihood in terms of profit generation, per I-1. Although not all projects that generate realized profits are scams, our observations in §4 show that all reported scams exhibit a positive realized profit. This indicator helps the heuristic identify and exclude projects with no profit.
The unrealized profit of the pool’s owner is another critic metric, especially when the SLID LP shows a significant growth in unrealized profit after the first month of deployment. The first-month unrealized profit represents the portion of profit (in the base tokens) belonging to the owner from the liquidity pool’s value after its initial month of deployment (detailed computing equation in Appendix C). In §4, we applied this unrealized profit metric for comparisons between SLID and rug pull scams, revealing a significant disparity between the average first-month unrealized profit ratios generated by the two groups (0.67% of the initial investment for rug pull and 156% of the initial investment for SLID pools). However, we set the threshold for the unrealized profit in this heuristic to 0 due to the uncertainty stemming from our current small initial ground truth dataset from RQ1, with the lowest computed unrealized profit for a SLID instance being $1.3e^{-14}$. This threshold addressed our I-2 of the scam likelihood pattern in generating unrealized profit.
We acknowledge that limiting the capturing range of unrealized profit to only within the first month might exclude potential SLID cases where the owner’s pool share develops afterward. Thus, a soft heuristic threshold of first-month positive unrealized profit allows a return with a lower bound on the ground truth dataset that only collects relevant cases with strong similarity to SLID discovered characteristics.
Owner activity validator.  SLID scams prioritize a gradual approach by multiple small-proportion profit-taking actions over an extended period (§5). We accorndingly built Owner Activity Validator in the following steps:
  • $\bullet$
    excludes the liquidity pools where the pool owners faithfully burn their initial (LP-) tokens after deploying the pool;
  • $\bullet$
    continue to detect the pools where their owners still getting all control over the initial (LP-) tokens;
  • $\bullet$
    observe/analyse the owner’s activities in exploiting profits.
We measure their activities through two metrics: (i) Profit-taking order count, denoted by $c$, accounts for the total number of DEX orders involved in profit extraction; and (ii) order impact, denoted by $I$, indicates the ratio of the $z^{th}$ profit-taking order size of value compared to the pool’s total value before the activity is executed. We define their thresholds as:
$c\geq{T_{count}},\,\text{where}\,I_{z}=\frac{y_{z}}{{x}_{t-1}}<T_{impact},\,\forall\ z\in\{1,2,\dots,c\}.$ (1)
where $T_{count}$ and $T_{impact}$ are the thresholds (§5) of owner’s profit-taking orders counts and maximum impact on the pool.
We select $T_{count}=5$ as the lower threshold for the profit-taking order count from the owner for their pool to be flagged as SLID. This choice is supported by our observations in §5: the SLID owner with the least profit-taking activities engaged in 5 profit-draining orders. Additionally, $I_{z}$ is computed as the order size, $y$, of the $z^{th}$ profit-taking order from the owner divided by the previous pool size ${x}_{t-1}$, where $t$ in this equation represents the order count of profit-taking order $z$.
To reflect the characteristic of small-proportion orders, we set the impact threshold of all owner’s profit-taking orders to be less than 95% ($T_{impact}<0.95$) of the total pool size. Our analysis in §5 revealed that 95% was the lowest profit-taking impact recorded for a rug pull instance. It also ensures we capture a broader range of scams, including potential SLID scams that have been unnoticed by rug pull detection heuristics. In comparison, for rug pull detection heuristics in [24, 37], only pools with a 0.99 profit-taking impact are flagged.
Additionally, we decide not to incorporate adaptivity for all numerical thresholds stems from the lack of ground truth in this new exploration of SLID . We focus on confirmed scams, including mislabeled rug pulls. We acknowledge that in real-world systems, attackers may adapt to the threshold and evade detection, but our model still lays a foundation.
[…]
6.2 Heuristic Evaluation (Small Scale)
We apply different datasets to evaluate our heuristic.
Accuracy on ground truth datasets. Before applying our heuristic to Ethereum liquidity pools (provided in §7), we first utilize our initial ground truth to evaluate the accuracy of the proposed rule-based heuristic. The rule-based heuristic has successfully flagged the 71 liquidity pools as SLID pools and correctly labeled the others as non-SLID scams. Evaluation results demonstrate that our heuristic achieved perfect accuracy (100%, experimental figures thus ignored) when tested on the RQ1 ground truth, resulting in no false positives. This lack of false positives in this evaluation is due to the limited ground truth available in this study. Since the heuristic is tailored directly from the ground truth, there is no other ground truth for false positive determination.
Heuristic performance on legitimate liquidity pools. To ensure that the heuristic generalizes well beyond our collected ground truth dataset and remains effective for identifying liquidity pools outside this dataset, we further validate its accuracy by testing it on external pools.
Given the absence of prior research or an established ground truth for SLID liquidity pool data, we evaluate the heuristic’s ’global accuracy’ by manually analyzing liquidity pools outside of our initial dataset. Specifically, we selected 100 randomly sampled legitimate liquidity pools deployed after 2022, consistent with the time threshold used in collecting rug pull reports (§5.1). We analyze the heuristic’s true negative and false positive rates (ability to correctly identify legitimate pools as non SLID without flagging them) to provide a robust measure of its performance.
With the heuristic, none of the 100 legitimate LPs were flagged as SLID scams. They all triggered the condition of burning initial LP tokens (defined in Owner Activity Validator). However, if we remove the LP burning condition from the heuristic, eight pools were flagged as SLID. These pools exhibited owner actions of gradually taking profits, similar to those seen in SLID scams. Notably, these actions were announced in advance through tokenomics, indicating a pre-declared intention rather than fraudulent behavior. This indicates that our heuristic still keeps 100% accuracy.
[…]
Findings of RQ2: The rule-based heuristic for SLID liquidity pool detection, with the validator for honeypot, profit and owner-action, demonstrates to be an effective detector. When tested against the initial ground truth and a dataset of 100 legitimate liquidity pools, it achieves 100% accuracy.
[…]
7 Detecting Ethereum’s SLID Scams (RQ3)
We conduct a large-scale scam detection analysis (Figure 5). We apply our rule-based heuristic (§6.1) on the datasets (including liquidity pools, and user activities) collected from six leading Ethereum DEXs (§7.17.2). In addition, we propose an ML-based approach (§7.3), designed to provide earlier detection of SLID scams. This is crucial for protecting investors from scams that unfold gradually over time.
Figure 5: Overall processes of our SLID detection.
7.1 Data Collection
We collected longitudinal Ethereum’s LP data spanning over 67 months. We preprocessed the data to extract relevant features. We utilized both heuristic-based and machine learning algorithms to classify the pools as SLID or non-SLID .
Liquidity pool information.  We first collected all the existing liquidity pools in Ethereum (denoted as PoolData) to be used for both detection methods. We utilized the well-known Geth client node [49] to retrieve all Ethereum block information, starting from block 6,627,000 (November 2, 2018, the date when the first liquidity pool was deployed) to block 20,081,867 (June 13, 2024, the time of writing this work). Then, we filtered only pools containing at least one well-established token (such as ETH, USDT, or USDC) as the pool’s legitimate base token. We avoid adding less-recognized and less-active pairs to the dataset.
As a result, we have gathered a total of 319,166 deployed liquidity pools across the six largest DEXs on Ethereum [50] (Table 1), inclduing Uniswap [33], Curve Finance [12], Balancer [51], SushiSwap [13], PancakeSwap [52], and BancorSwap [53]. We extracted key features (Appendix G) for each harvested liquidity pool, including creation timestamps, token addresses involved in the pool, owner addresses, and the initial pool value in USD, using Chainalysis API [54].
Table 1: Collected data on Ethereum.
Liquidity Pools DEX orders SC Features
DEXs No. of pools Order type No. of Txs No. of SCs
UniSwap 312,299 302,675
SushiSwap 2,898 Deposit 12,207,714
Balancer 2,819 Withdraw 5,671,266
Curve 294 Buy 505,364,145
PancakeSwap 443 Sell 394,416,456
BancorSwap 413
Total 319,166 917,659,580 302,675
  • Abbreviation: Tx: transaction; SC: smart contract; No.: number.
DEX orders. Exchange activities and profit patterns within PoolData are essential for computing the profit-related metrics and profit-generating metrics (§6.1). We collected all types of exchange activities (via smart contract events) of six DEXs. Those activities include liquidity deposit, withdraw, and trade (buy/sell) between two tokens in each pool. Considering the data collection process operates independently of PoolData extraction, we specifically filtered the orders to include only those relevant to the liquidity pools in PoolData.
We collected over 917M DEX activities across 319,166 liquidity pools (Table 1). It accounts for 82.35% of the total Ethereum trading volume historically [55]. The resulting dataset, referred to as DEXData, contains rows where each entry represents an activity by either the pool owner or an investor within a specific liquidity pool. The dataset includes activities recorded up to July 13, 2024, where all pools in PoolData have at least one month of DEX activity history.
Smart contract security features.  Token security-related features (such as buy/sell tax, trading lock, etc.) from the paired token’s smart contract are critical for applying the heuristic’s Honeypot Validator. We utilized the security API provided by GoPlus [43] and De.Fi [39] to collect the token’s security characteristics indicative of honeypot behavior, including buy/sell tax, sell restriction, and more.
We extracted security features from 302,675 distinct tokens (base and paired) across 319,166 Ethereum liquidity pools, noting that some tokens are associated with multiple pools. These features were organized into a dataset, referred to as SecurityData, where each row represents a distinct token along with its complete set of security-related features.
7.2 Applying Heuristic for SLID Detection
We described how we apply the heuristic to detect Ethereum’s SLID scams. We began with data preprocessing and extracted the required features for all the liquidity pools. We then conducted an in-depth analysis of the flagged SLID pools across multiple dimensions, including vitality, profit, and trend.
Data preprocessing.  We preprocessed the three datasets (PoolData, DEXData, and SecurityData) to enhance the feature set and prepare for the heuristic detection phase.
For each liquidity pool instance in PoolData, we extracted all related DEX activities from DEXData to compile its complete activity history. Profit-related and owner-activity metrics were then computed (equations in Appendix C) and added as new features to the PoolData dataset. These enriched metrics serve as key inputs for the flagging process. Then, smart contract security features from SecurityData were mapped to their corresponding liquidity pools in PoolData.
Our preprocessing resulted in an updated dataset, called ExtendedPoolData, containing a comprehensive range of features (Table 4) required for heuristic detection.
SLID detection for Ethereum. We applied the proposed rule-based heuristic to all instances in the ExtendedPoolData dataset. The new dataset now contains all required features for processing with the heuristic’s validators.
  • $\bullet$
    (Quick access to our dataset) Out of 319,166 selected liquidity pools, we flagged 3,117 instances of SLID scams. For evaluation, we randomly selected 100 SLID pools for manual observation, which revealed no sign of false positives. To validate the novelty of this scam tactic, we tested the randomly selected pools using existing honeypot and rug pull heuristics, none of which detected the scam.
(i) Vitality analysis of SLID pools. Figure 8 depicts the SLID liquidity pool age and the remaining active pool numbers in each age group. Of the 3,117 SLID liquidity pools, 70.8% (2,207/3,117) remained active after the first month, and 53.3% (1,660/3,117) were still operational three months after their deployment. Five pools reached the longest recorded age of four years. Among the 196 pools that were active for over three years, 28 are still alive today. The performance is exceptional compared to other scams (rug pull and honeypot). Over their lifetime, SLID pools attracted a total of 1.49M activities on DEXs, with a cumulative volume of \$4.49B.
We concluded that the SLID scams exhibit the characteristics of a long pool lifespan, with remarkably extended periods of activity and significant user engagement. This result can be further supported by their total trading volume.
Figure 6: Lifetime analysis of captured SLID liquidity pools.
Figure 7: Daily realized profit drained and scammer’s profit-taking activity count.
Figure 8: Daily volume (in USD) and activity count in SLID scam liquidity pools.
(ii) Profit analysis of SLID pools. We analysed the financial earnings of scammers. Across all flagged SLID liquidity pools, the scammers have lured 167,632 victims, generating \$79.1M in profit through a gradual profit-taking strategy. Additionally, unrealized profit assets worth \$24.1M remain under the control of scammers in active pools. On average, scammers achieved a profit of 6.08 times their initial investment. The top liquidity pools generating the highest profit are described in Table 2. The 3,117 SLID scams were initiated by 2,951 distinct addresses, with 90 addresses involved in more than one scam and one address deploying a peak of 23 SLID scams (details in Table 3).
We further investigated the scammers’ gradual strategy by analyzing their profit-taking patterns over time from the first day of pool deployment. Figure 8 illustrates the total number of profit-taking actions and the corresponding realized profit by days from each pool’s respective deployment date. We observed that most profit-taking actions occur in the first 100 days of the SLID liquidity pools, with 35.71% (31,912/89,364) of sell or withdraw actions executed during this period. While the number of profit-taking actions decreases over time, the total realized profit remains relatively stable throughout the evaluation period, ranging between \$1K and \$100K, with a few outliers.
We concluded that the scammers, who leveraged the SLID strategy, have the alarming ability to continue generating profits even during later stages with fewer profit-taking actions.
(iii) Trend analysis of SLID pools. We analysed the trend of collected pools. Figure 9 illustrates a steady increase in the number of SLID pools created annually, with the exception of 2024, where the lower number is due to the limited capture period (6 months). Meanwhile, Figure 8 illustrates the total activities and volume associated with SLID scams. We describe the development of SLID scams as follows.
  • $\bullet$
    Early development phase (2018 - 2020): During this period, a small number of SLID pools were created, exhibiting low volume and minimal activity. This could be attributed to the limited presence of DEXs and the lack of user familiarity with decentralized activities during this time [56]. However, there was a gradual increase in the development of SLID LPs, with recorded increments in user engagement.
  • $\bullet$
    Exploding phase (Mid-2020 - 2023): Following the rapid growth of DeFi, the market experienced a significant surge in SLID scams. The number of SLID liquidity pools created in 2020 was over 10 times greater than in 2019, accompanied by an explosion in trading volume and activity within these pools, with volumes ranging between \$10K and \$100M. This phase also saw record-breaking performances for SLID scams, with an all-time peak total volume of \$85.7M on March 8, 2021, and peak DEX activity counts of 17,586 on December 20, 2022, across all SLID LPs.
  • $\bullet$
    Stabilized phase (2023 - present): The statistics for SLID liquidity pools maintained high levels of stability, with daily activity counts ranging from 100 to 1000 and daily volumes ranging from \$10K to \$10M. During this period, the average daily volume was \$608,159, and there were 266.67 DEX activities per day across SLID pools. Although not exhibiting the same rapid growth as the previous one, the scam tactic remains attractive to scammers, as evidenced by the creation of 1,107 new pools. This continued appeal is likely due to the lack of prior recognition.
We conclude that there is a clear upward trend in the initialization of SLID scams over the years. The SLID strategy has been increasingly adopted by malicious actors.
7.3 ML Approach for Early SLID Warning
We take a step further. Our heuristic can only flag SLID pools only after profit-taking actions that trigger the heuristic’s validators. However, investors already lost their money.
We thus propose a proactive ML-based approach to raise early warnings and prevent users from engaging at the pool’s early stages. We utilized machine learning methods to enable fast detection. In particular, we extracted features from the liquidity pool in progressively shorter time windows to identify a reasonable timeframe that offers high prediction accuracy.
2018201920202021202220232024005005005005001,00010001{,}0001 , 0006464646452525252585585585585446446446446863863863863979979979979128128128128Year (p.s. 2024 up to June)SLID LPs
Figure 9: Statistic of the created SLID LPs over years.
Data preprocessing and feature extraction.  During our heuristic detection phase, we labeled all the collected liquidity pools. We still utilized these labels as ground truth for training the ML algorithms. It contains 3,117 flagged SLID pools and 316,049 non-SLID pools in the ExtendedPoolData dataset.
For early detection, we constrained the liquidity data (i.e., DEX activities) used for feature extraction to the first $d$ days after the deployment of pools in ExtendedPoolData.
During the preprocessing phase, we extracted features for each ground truth liquidity pool within the first $d$ days of its creation and appended these features to the corresponding liquidity pool records in ExtendedPoolData.
We then utilized the resulting dataset, called $d^{th}$-day, to train and evaluate ML algorithms. We varied the $d$ value, training the algorithms separately for each duration. We determine the shortest time frame by systematically reducing $d$, which balances high accuracy in detecting SLID scams with early warning capabilities. For pools with a lifetime shorter than $d$, features were extracted based on all available DEX activities from deployment to their inactive state.
To build an effective classifier within the specified time constraints, we extracted 57 features from DEXData, capturing diverse patterns of DEX activities. These features were categorized into four groups: Owner Activity Features (OAF), User Activity Features (UAF), Liquidity Pool Features (LPF), and Profit Features (PF) (details in Appendix G).
ML algorithms. We framed the early detection problem as a binary classification task, focusing on models that deliver high accuracy while maintaining low complexity. Specifically, the model selection criteria emphasized efficiency, model complexity, computational cost, and accuracy, requiring a careful balance among these factors.
Based on these considerations, we focused on three ensemble machine learning models: Random Forest [57], XGBoost [58], and Logistic Regression [59]. We exclude several alternative approaches, such as graph-based or network-based strategies, due to their high complexity and computational demands. Moreover, they are unsuitable for the scenarios we address [60] (more rationales in Appendix E).
Early detection. We measured their effectiveness across various $d$ values using metrics of accuracy, precision, and recall.
To address the imbalance problem, we applied class weighting [61] and utilized GridSearchCV from sckikit-learn [62] for hyperparameter tuning for each $d$-algorithm combination. To demonstrate the detection efficiency in a shorter time period, we conducted an experiment using the heuristic on the 3,117 pools (pseudocode provided in Algorithm 1).
We selected the initial value $d$ as 267 days because, within this time frame, all four detectors remain to demonstrate optimal performance in SLID detection.
Detection results (Figure LABEL:fig:detection-metrics). As the time range was reduced, the performance of recall of the heuristic began to decline due to the limited number of order activities, which caused the heuristic validators to fail to trigger some detections. Given that the true negative count in the ground truth (316,049) is substantial compared to the true positive (3,117), both accuracy and precision of the heuristic remained unchanged when $d$ decreased. In contrast, the three ML algorithms maintained stable and good performances with decreasing time ranges. However, Random Forest performs consistently better than the other two ML algorithms in most $d$.
Random Forest was the last to maintain an accuracy greater than 93% down to $d$ = 56, whereas Logistic Regression and XGBoost saw their accuracy drop to 93% when using data within a time range of 58 days from the liquidity pool’s deployment date. Starting from $d$ = 56, the accuracy of all ML algorithms fell to 93%, marking the threshold at which all algorithms lost their high performance observed previously.
Overall, the three ML algorithms demonstrated detection capabilities that were 4.77 times faster than the proposed heuristic, achieving high accuracy with just 57 days of data compared to the heuristic’s requirement of 268 days.
[…]
8 Discussions
8.1 SLID Scam Mitigation
SLID remains concealed and challenging for users to detect quickly. However, we stress several key points worth noting.
  • $\bullet$
    Verify token purpose and cross-check the claims. Ensure the token aligns with its intended purpose (utility, governance, investment). Any discrepancies between official claims and token behavior may signal a scam.
  • $\bullet$
    Evaluate LP tokens. Confirm that initial LP tokens are burned or securely locked using trustless mechanisms. Tools like Etherscan, DEXTools, and Chainalysis can help with verification. Unburned/unlocked tokens are red flags.
  • $\bullet$
    Assess on-chain data. Prioritize transparency and legitimacy by thoroughly evaluating on-chain activities. Regularly monitor the token to detect any signs of the owner gradually draining the liquidity pool.
8.2 SLID Limitations and Our Plan
Exceptional ambiguous cases. Initial LP token burning is an industry standard for ensuring pool transparency and preventing the owner from controlling 100% of the pool’s funds. However, this study does not cover cases where liquidity pools are created without burning LP tokens, and the owner takes profits only a few times, appearing legitimate. Such cases may initially resemble SLID patterns but later gain legitimacy as the token becomes popular (e.g., with meme coins or community-driven tokens).
These pools operate on the grey zone of the boundary between fraudulent activity and legitimate trading strategies while containing all fraudulent traits (i.e., the owner controls the pools with all investor’s money, can take them all anytime, and can do so multiple times through small profit-taking activities) but gain community approval. Those ambiguous cases make them challenging for us to classify as fraudulent confidently. Although the evaluation didn’t show the result of any involved legitimate liquidity pool being falsely labeled as SLID, we acknowledge this limitation that the heuristics (and ML models) might have limited ability in those cases and might result in false positives. Our future work aims to refine and update current heuristics to account for these ambiguous cases, potentially involving expanding features and community sentiment in heuristic labeling.
Advanced manipulations. Although our designed heuristic includes evaluations of the initial ground truth and legitimate tokens, we cannot yet address complex manipulations. We provide two example scenarios: (i) the liquidity pool conducts initial token burning for transparency and still attempts to commit with SLID actions; (ii) the pool owner does not conduct the action themselves but transfers LP tokens/inflated paired token amounts to another/others account(s) to execute a SLID strategy. We acknowledge that the true scope SLID may extend beyond what we detect, and we treat the present results as an upper bound and recognize the need for further refinements in the future.
8.3 Real-world Integration
SLID monitoring and early warning services. Our solution can be integrated into DEXs or crypto-related services as a LP monitoring service, a cross-checking mechanism, or a new security standard/requirement. A key requirement is the extraction of features, which can be directly derived from the token’s transaction history to evaluate LPs deployed on public blockchains, in particular, Ethereum. This service provides real-time warnings to investors about any fraudulent activities in existing or newly deployed tokens.
However, there are challenges. The dynamically required features for the ML models (e.g., volume and owner activity counts, which continuously change) demand significant computational resources to maintain and update. If a constant monitoring service is to be provided based on our solution, the model would require substantial resources to track and process these evolving features from the targeted pool. This limitation creates a barrier, making it nearly impossible to build an SLID detection service that can constantly update features and monitor all existing liquidity pools.
To date, the service can focus on a targeted set of suspect liquidity pools, enabling more effective resource allocation while still maintaining early detection.
Wallet credibility evaluation. Most payment provider services accept cryptocurrency transactions only from legitimate and well-established tokens. For instance, PayPal accepts cryptocurrency payments from decentralized wallets using Bitcoin and a few other valid tokens [63]). Newly created tokens are rarely considered by legitimate payment providers.
Integrating our solution with payment services enhances risk awareness in crypto-transactions, even if the SLID token is not yet listed. By identifying wallets linked to SLID creation as fraudulent and analyzing their transaction histories, our solution detects SLID activity and assesses fraud risks. This enables providers to flag fraudulent wallets, even when handling legitimate tokens like Bitcoin.
The limitation is the uncertainty of the required time to process transaction history to extract features. The computation time varies depending on the length of a pool’s historical transactions. One potential workaround is to provide offline detection. However, the continuous updating of pool transactions and their features presents a challenge.
9 Conclusion
This paper introduces SLID scams as a stealthy and financially devastating threat to the DeFi ecosystem. By combining heuristic-based detection with machine learning, we identified 3,117 SLID pools, exposing \$103.2M in cumulative losses. Our ML model achieves early detection, enabling timely intervention to protect investors. We have responsibly disclosed our findings to the affected DEXs. We hope this study raises community awareness to prevent such gradual, daily losses. “Slow” does not mean safe; it can silently drain your wallet.
Appendix A Additional Related Work
Liquidity pool, DEX and DeFi. Over the past five years, the rise of DeFi and DEXs has attracted millions of users and billions in inflows, leading to numerous research studies [64, 65, 66, 56, 9, 67, 68, 69]. Hendrik et al. [64] examined the opportunities and challenges for DeFi and its products in this mass adoption era. Fan et al. [66] conducted a comprehensive study on cryptocurrency trading activities and methods for measuring market valuation. Jianlei et al. [56] investigated the DEXs roles and their impact on the development of the DeFi ecosystem. Saleh et al. [67] analyzed the effects of liquidity on Uniswap for liquidity providers and conducted estimations of fees and impermanent loss. Similarly, Lioba et al. [68] developed a theoretical model to analyze the risks and returns faced by liquidity providers under various circumstances.
Cryptocurrency scam detection. Various research efforts have focused on analyzing DEX scams, such as rug pull [24, 37, 70, 71] and honeypot [25, 48, 72, 73]. For instance, Bruno et al. [24] utilized ML algorithms to identify rug pull and scam-likely tokens. Cernera et al. [37] conducted a longitudinal analysis of the lifetime of rug pull liquidity pools on Ethereum and Binance Smart Chain. Rundong et al. [48] introduced various types of honeypot traps on DEXs along with their respective detection methods. Similarly, Christof et al. [25] conducted a comprehensive honeypot analysis with a proposed symbolic-execution-based solution. In addition, the scams originating from traditional finance, such as pump-and-dump schemes [74, 75] and Ponzi schemes [44, 76], have also been studied to provide warnings to investors. Several studies [77, 78, 79, 80] have suggested general fraud detection models that identify anomalous transactions or users through graph neural networks. However, among all the studies in this field, the SLID scam has yet to be discovered.
Appendix B Pool Owner Financial Guarantee in SLID
Let the liquidity pool’s initial deposit (by the owner) be $x$ units of the paired token and $y$ units of the base token. From the formula of the AMM model from Uniswap [34], the relationship between these tokens in the pool is reflected by:
$x\times y=k$ (2)
where $k$ is a constant determined by the initial liquidity and remains unchanged. With the mentioned setup, the pool’s owner initially monopolizes the paired token supply, which means no paired tokens $x$ exist outside the pool before the pool’s deployment. Any investor/user that is investing/interacting with the pool must first buy the paired token from the pool by swapping it with base tokens. Let $\Delta x$ represent the total amount of the paired token bought by users or distributed as incentives (for users that deposit a pair of tokens for liquidity providers). The formula will be rewritten to represent the pool’s state any moment after deployment:
$\begin{split}(x+\Delta x)\times(y+\Delta y)=k\\ x^{\prime}\times y^{\prime}=k\end{split}$ (3)
where $\Delta y$ is the change of base token amount in the pool due to the change of paired token, and x’/y’ is the amount of paired/base token in the pool after the change compared to the initial deposit amount (from the owner). Under the AMM model, the constant $k$ must not change.
In the worst scenario where no investor interacts with the pool ($\Delta x=0$) or all investors sell back their paired token ($=\Delta x$, since the monopoly of paired tokens from the owner made the limit of paired token that investor possess totally at any time cannot be larger than $\Delta x$) to the pool, the base token possessed by the owner afterward is computed by:
$y^{\prime}=\frac{k}{x^{\prime}}=\frac{x\times y}{x+\Delta x-\Delta x}=y$ (4)
This proves that in the worst-case scenario, where the pool has no investor or all investors that previously bought (and use the bought paired token to deposit for incentives) sell back all their received paired tokens to the pool, the base token balance controlled by the owner will still equal to $y$ which is the initial investment of the owner to the pool as initial liquidity. Hence, with this setup, the pool owner is guaranteed to maintain their original investment of base tokens $y$, which contain real value, under any sequence of investor interactions after the pool’s deployment.
Appendix C Computing Equation
Realized profit computing equation. To compute the realized profit, we let $m$, $n$, $o$, $p$ represent the total number of owners’ sell, withdraw, buy and deposit orders to their pool, respectively. The realized profit can be defined as follows:
$\begin{split}\textit{Profit\_realized}=&\left(\sum_{i=1}^{m}Sell_{i}+\sum_{j=1}^{n}Withdraw_{j}\right)\\ &-\left(\sum_{k=1}^{o}Buy_{k}+\sum_{l=1}^{p}Deposit_{l}+\text{GAS}\right)\end{split}$ (5)
where the owner’s investment activity (which adds assets in base tokes to the pool) includes $Buy_{k}$ and $Deposit_{l}$, representing the asset value (in USD) added to their liquidity pool during the $k^{\text{th}}$ buy and $l^{\text{th}}$ deposit orders, respectively. In contrast, the owner’s realized return consists of $Sell_{i}$ and $Withdraw_{j}$, which are the asset values in base token (in USD) that the owner removed from the pool during the $i^{\text{th}}$ sell and $j^{\text{th}}$ withdraw orders, respectively. The metric also accounts for the total network gas fee (GAS) used for pool deployment and all DEX activities performed by the owner. Since the equation focused on computing realized profit, all of the variables are only around the owner’s balance change of the base token (which holds the real value) through their activity in the pool (which also involves paired token as well).
Unrealized profit computing equation.  To compute this unrealized profit, we begin with the pool’s value (in USD), denoted as $x$, in the base token, and $owner\_share$, representing the pool owner’s portion of the pool’s asset. During the pool deployment phase, $x$ will be the owner’s initial deposit value: $x=x_{0}=Deposit_{0}$ and $owner\_share$ is 1, as they initially own 100% of the pool. For each subsequent $t^{th}$ DEX order made to the liquidity pool, the value of $x$ is updated as follows:
$x=x_{t}=x_{0}+\sum_{p=1}^{t}y_{p}$ (6)
where $y_{p}$ represents the value of the base token amount (USD) for the $p^{th}$ DEX order. If the order is a buy or deposit (indicating a positive $y$), it signifies that $y$ USD worth of base token has been added to the pool through the order. Conversely, sell and withdraw orders indicate a decrease in value (indicating a negative $y$) due to the outflow of base tokens. Furthermore, for each “deposit” and “withdraw” order, the pool proportions shared amongst the liquidity providers change by,
$owner\_share_{t}=\begin{cases}\begin{array}[]{l}owner\_share_{t-1}\times\frac{x_{t-1}}{x_{t}}+\frac{y_{t}}{x_{t}}\\ \text{(if $t^{\text{th}}$ is a deposit/withdraw activity}\\ \text{from the owner),}\end{array}\\ \begin{array}[]{l}owner\_share_{t-1}\times\frac{x_{t-1}}{x_{t}}\\ \text{(if $t^{\text{th}}$ is a deposit/withdraw activity}\\ \text{from other investors).}\\ \end{array}\end{cases}$ (7)
Eq. 7 indicate that ${owner\_share}_{t}$ increases if the owner deposits more to the pool, and conversely, it decreases if the owner executes withdraw orders. Similarly, the owner’s share will decrease if a user deposits into the pool and will increase if a user withdraws from it. The mechanism behind the owner pool’s share in Eq.7 is explained in more detail in D. Therefore, the liquidity pool’s owner’s first-month unrealized return is defined as follows:
$\begin{split}\textit{Profit\_unrealized}_{1m}=\textit{Return\_unrealized}_{1m}\\ ={x}_{t}\times{owner\_share}_{t},\\ \text{where $\textit{Profit\_realized}_{1m}\geq 0$}\end{split}$ (8)
where $t$ in this equation represents the order number executed in the pool on the DEX since its deployment to the end of the first month. If the liquidity pool has a positive realized profit threshold, indicating that it has reached the break-even point and already generated profit, we can consider the unrealized return as unrealized profit since it represents a profit beyond the realized gains.
Appendix D Owner’s Pool Share
We explain how the owner’s liquidity share is updated or changed after a deposit or withdrawal in a DEX liquidity pool. We denote the share of owners at the $t^{th}$ order as $\mathcal{S}_{t}$. We present two scenarios to illustrate this process.
  • $\bullet$
    Deposit/withdraw excluding the onwer.  When a deposit or withdrawal order is executed in the pool and the sender is not the owner, the owner’s share is updated as follows:
    $\mathcal{S}_{t}=\mathcal{S}_{t-1}\times\frac{x_{t-1}}{x_{t}}.$
    The owner’s liquidity pool share before the $t^{th}$ order is calculated by multiplying their previous share $\mathcal{S}_{t-1}$ by the pool value before the order $x_{t-1}$. After the $t^{th}$ order, the updated share is determined by dividing the owner’s liquidity amount by the new pool value $x_{t}$.
  • $\bullet$
    Deposit/withdraw by the owner. When the pool owner executes the $t^{th}$ order with size $y_{t}$, we compute the new owner share as follows:
    $\mathcal{S}_{t}=\mathcal{S}_{t-1}\times\frac{x_{t-1}}{x_{t}}+\frac{y_{t}}{x_{t}}.$
    Whether the $t^{th}$ order involves a deposit or withdrawal, the ratio $\frac{y_{t}}{x_{t}}$ represents the percentage of the new liquidity part contributed by the owner relative to the pool’s value. This ratio, reflecting either an addition or subtraction, merges with the owner’s share prior to the $t^{th}$ order and updates their stake accordingly. If the $t^{th}$ order is a deposit by the owner, the ratio $\frac{y_{t}}{x_{t}}$ yields a positive value, as the pool value $x_{t}$ is inherently positive, thereby increasing the owner’s share post-update. Conversely, in the case of withdrawals, this ratio becomes negative, diminishing the owner’s share.
Appendix E Rationale for Seletected MLs
We chose Random Forest, XGBoost, and Logistic Regression because they offer:
  • $\bullet$
    efficiency for real-time detection: the models are computationally efficient once trained. They can process new transactions quickly, ideal for (near-)real-time scenarios. They also scale well to high-throughput environments, such as mempool monitoring or transactional analysis, where rapid predictions on incoming data are critical.
  • $\bullet$
    compatibility with feature-based data: They excel in structured, tabular data environments, leveraging numerical and categorical features without requiring complex data transformations, unlike graph-based approaches (require transforming data into relational structures).
  • $\bullet$
    balance of complexity and accuracy: Random Forest and XGBoost capture non-linear patterns and interactions, while Logistic Regression provides a lightweight yet robust alternative for simpler cases, ensuring high performance with manageable complexity.
  • $\bullet$
    transparency and interpretability: These models provide clear insights into feature importance (Random Forest) or coefficients (Logistic Regression), aiding decision-making and enabling user trust in the detection process.
Appendix F Additional Statistics
Table 2 presents the top 5 LPs with the highest profits. Table 3 shows the scammers with the most SLID LPs and total profits.
Table 2: LPs with the highest total profit generated
Pool address Paired Token Profit (USD)
0x32Ce…cb81b CORE \$4,261,805
0xBE5…e0732 OptionRoom \$3,673,710
0x092…DB485 Public Mint \$2,470,875
0x8d9…c7f5F Finxflo \$1,174,583
0x9dc…198Fd Bondly Token \$986,684
Table 3: SLID scammers that deployed most SLID LPs
Scammer address SLID scam created Profit (USD)
0x3DC…1f887 23 \$17,307,795
0x4bb…ca0a0 15 \$395,054
0x285…409c1 9 \$298,459
0x521…dCDD2 7 \$2,064
0x06c…AD87C 6 \$4,170
Appendix G Data Collection and Features
Collected data. We report the statistics of collected data (Table 1) used in the data collection process (cf. §7.1).
Gathered features.  We show features (Table 4’s upper-side) from collected LPs and DEX order events from PoolData. We gather token security features for honeypot determination.
Extracted features in ML detectors.  We show 57 features (Table 4’s bottom-side) extracted from DEXData within a given time window $d$ for the training and testing activities of the ML detector. The features are categorized as:
  • $\bullet$
    Owner activities features (OAF). These features reflect the owner’s interaction behavior with their LPs, including their activity statistics within the time range of $d$.
  • $\bullet$
    User activities features (UAF). This category includes metrics that capture user activities during this period, providing information about both the users and their activities.
  • $\bullet$
    Liquidity pool features (LPF). These illustrate the characteristics of LPs during the specified period, encompassing aspects of the pool’s lifetime, volume, and asset value.
  • $\bullet$
    Profit features (PF). Features in this category indicate the profit patterns generated by the liquidity pool for its owner, covering a wide range of metrics, including return, profit, and return on investment.
Appendix H Pseudocode
Our main program consists of three functions (Algorithm 2), each focusing on an aspect of the liquidity pools’ operations.
Data: DEXData
1
2 ML_approach (SLID_pools, DEXData) :
3       foreach d in 1 to 300 do
4             ExtendedPoolData_[d] $\leftarrow$ Extract features from DEXData in transaction range between day 1 to day d
5             results[d][heuristic] $\leftarrow$ evaluation metrics (accuracy, precision, recall) when applying heuristic on ExtendedPoolData_[d]
6             /* get all SLID pools */
7             foreach model in [RandomForest, LogRegression, XGBoost] do
8                   trained_model $\leftarrow$ Train model using ExtendedPoolData_[d]
9                   evaluation $\leftarrow$ Evaluate trained_model on validation/test set
10                   results[d][model] $\leftarrow$ evaluation metrics (accuracy, precision, recall)
11
12             end foreach
13
14       end foreach
15
16
Algorithm 1 ML approach for early SLID warning
The first function (LP_Age_Analysis) determines the longevity and activity status of each pool. It retrieves transactions associated with each pool and calculates the pool’s age by comparing the earliest and latest transaction timestamps. This data helps assess how long each pool has been operational. The function also tracks whether pools are still active by checking for recent transactions.
In the second function (LP_Profit_Analysis), the focus shifts to the financial activities of the pool owners. This part of the algorithm targets transactions where the pool owners are selling or withdrawing assets. By analyzing these transactions, the function calculates the total volume of assets moved, providing insights into the profit-taking behaviors of pool owners.
The final function (LP_Trend_Analysis) examines the engagement of other users within the pools, excluding the owners. It analyzes transactions made by users to understand the general activity level and trends within each pool. This function helps to highlight user participation and can indicate the popularity of the pool based on user activity.
We further present pseudocode for our enhance method by using machine learning (Algorithm 1).
Over 300 days, our algorithm extracts daily transaction features to evaluate and compare heuristic and machine learning models (RandomForest, Logistic Regression, XGBoost) based on accuracy, precision, and recall.
1
Data: ExtendedPoolData, DEXData
2
3SLID_pools $\leftarrow$ ExtendedPoolData.label == ’SLID’
4 /* get all SLID pools */
5
6Token_Pool_Age_Analysis (SLID_pools, DEXData)
7LP_Profit_Analysis (SLID_pools, DEXData)
8LP_Trend_Analysis (SLID_pools, DEXData)
9 LP_Age_Analysis (SLID_pools, DEXData) :
10       foreach pool in SLID_pools do
11             pool_txs $\leftarrow$ all entries in DEXData where DEXData.pool_address == SLID_pools.pool_address
12             /* get all transaction of the pool */
[…]
14             /* pool age calculation */
[…]
16             /* add to list for graphing */
[…]
18             /* record alive pools of age groups */
19       end foreach
20
21
22
23 LP_Profit_Analysis (SLID_pools, DEXData) :
24       foreach pool in SLID_pools do
25             pool_txs $\leftarrow$ all entries in DEXData where DEXData.pool_address == SLID_pools.pool_address & DEXData.sender == SLID_pools.pool_owner & DEXData.category == sell/withdraw & DEXData.token_address == SLID_pools.base_address
26             /* get pool’s owner profit_taking activity */
27
28            foreach tx in pool_txs do
29                   volume[transaction.timestamp] $\leftarrow$ $tx[amount\_token]\times tx[token\_price]$
30
[…]
32
33             end foreach
34
35       end foreach
36
37
38 LP_Trend_Analysis (SLID_pools, DEXData) :
39       foreach pool in SLID_pools do
40             pool_txs $\leftarrow$ all entries in DEXData where DEXData.pool_address == SLID_pools.pool_address & DEXData.sender != SLID_pools.pool_owner & DEXData.token_address == SLID_pools.base_address
41             /* get all user activity of the pool */
42
43            foreach tx in pool_txs do
44                   volume[transaction.timestamp] $\leftarrow$ $tx[amount\_token]\times tx[token\_price]$
45
[…]
47
48             end foreach
49
50       end foreach
51
52
Algorithm 2 Liquidity pool analysis
Ethical Considerations
Our research focuses on uncovering and mitigating the slow liquidity drain (SLID) scam within the DeFi ecosystem. Given the sensitive nature of this study, we adhered to established ethical guidelines to ensure responsible data handling, minimize harm, and promote fairness across all stakeholders.
Privacy considerations. We focused our research on the behavior pattern of the owner that creates SLID liquidity pool and their activities. All examined data are publicly available by the Ethereum network, with all addresses anonymized by only keeping the first/last 5 characters of each, and generated mapping values for similarities determination (of determining transactions involving the same tokens/pools/users) in analysis tasks. We have no intention of deanonymizing any of the involved addresses for privacy reasons.
Respect for persons. We prioritized user privacy throughout this research. All data analyzed was sourced from publicly available blockchain records, and no personally identifiable information (PII) was collected or processed. While blockchain data is inherently pseudonymous (also claimed before), we ensured that our findings respected user anonymity. Additionally, any insights derived from community-reported rug pulls were aggregated to avoid targeting individual users.
Beneficence. To maximize benefits while minimizing potential harm, we aimed to alert the DeFi community to the risks posed by SLID scams. Our research introduces early detection mechanisms, equipping users with tools to make informed decisions and reduce financial losses.
To mitigate the risk of misuse, we avoided publishing details that could enable malicious actors to refine or replicate SLID scams. Instead, the study emphasizes preventive strategies and raises awareness among users and researchers.
Justice. The benefits of this research are distributed equitably across the DeFi community. By identifying a previously underexplored scam type, we empower both users and platform operators to take proactive measures. Our findings are not intended to single out specific entities but rather to foster collective improvements in security and trust.
Respect for law and public interest. This study was conducted in compliance with relevant regulations surrounding data privacy and blockchain technology. All activities were limited to publicly accessible data, and no vulnerabilities were exploited during the research. Instead, we aim to enhance user protection through transparency and collaboration with the broader DeFi community.
Proactive community engagement. Throughout this study, we engaged with community stakeholders, including affected users and DEX platforms, to validate our findings and refine detection strategies. By fostering open dialogue, we sought to align our objectives with the broader goal of improving DeFi security while addressing community concerns.
By adhering to these ethical principles, our research aims to strike a balance between innovation and responsibility, contributing to a safer and more resilient DeFi ecosystem.
Open Science
To promote transparency and reproducibility in research, we will share relevant details, datasets, and code used within this study available once the paper is fully accepted/finalized.
Table 4: Summerized features for datasets.
Feature Description Feature Description
Features collected for LPs, DEX activitivies and honeypot deterniniation
LP $Address_{pool}$ Ethereum address of LPs $Address_{base}$ Ethereum address of base token in LPs
$Address_{paired}$ Ethereum address of paired token in LPs $Address_{owner}$ Ethereum address of the pool owner
$CreatedTime_{token}$ Created time of the paired token $CreatedTime_{pool}$ Deployed time of LP
$DEX$ DEX where the liquidity pool deployed $Name$ Paired token’s name
DEX activities $Block$ Block number that the activity occurred $Timestamp$ Timestamp that the activity occured
$Hash$ Transaction hash of the activity $Category$ Activities (buy/sell/deposit/withdraw)
$Address_{pool}$ Ethereum address of LP $Address_{sender}$ Ethereum address of the sender
$Address_{base}$ Ethereum address of base token in LPs $Address_{paired}$ Ethereum address of paired token in LPs
$x_{paired}$ Pool balance of paired token after activities $y_{paired}$ Token amount of paired token
$x_{base}$ Pool balance of base token after activities $y_{base}$ Token amount of base token
$Price_{paired}$ Price of paired token recorded in the block $Price_{base}$ Price of base token recorded in the block
Honeypot $Tax_{buy}$ Buy tax of the token $Tax_{modify}$ Owner’s capability to change buy/sell
$Tax_{sell}$ Sell tax of the token $Buyable$ User’s capability to buy tokens
$SellAll$ User’s capability to sell all tokens $BalanceChange$ Owner’s capablity to change user’s balance
$TradingCooldown$ Time limit for tokens between activities $TradingPausable$ Owner’s capbility of pausing trading
$AntiWhale$ token activities with high amounts
Extracted features of OAF, UAF, PF and LPF for ML detector
OAF $Owner_{dep}$ Total owner deposit activity $Owner_{buy}$ Total owner buy activity
$Owner_{with}$ Total owner withdraw activity $Owner_{sell}$ Total owner sell activity
UAF $User_{dep}$ Total users deposit activity $User_{with}$ Total users withdraw activity
$User_{buy}$ Total users buy activity $User_{sell}$ Total users sell activity
$User_{count}$ Total pools user count
$User_{countfirst}$ User count first day of pool deployment $User_{counthigh}$ User count of day with most pool users
$User_{countlast}$ User count last day of pool deployment $User_{countlow}$ User count of day with least pool users
$RUser_{firstonhigh}$ The ratio of $User_{countfirst}$ and $User_{counthigh}$ $RUser_{lastonlow}$ The ratio of $User_{countlast}$ and $User_{countlow}$
$RUser_{firstonlow}$ The ratio of $User_{countfirst}$ and $User_{countlow}$ $RUser_{firstonlast}$ The ratio of $User_{countfirst}$ and $User_{countlast}$
$RUser_{lastonhigh}$ The ratio of $User_{countlast}$ and $User_{counthigh}$ $RUser_{lowonhigh}$ The ratio of $User_{countlow}$ and $User_{counthigh}$
PF $Owner_{i}$ Total owner investment (USD) $Owner_{u}$ Total owner unrealized return (USD)
$Owner_{r}$ Total owner realized return (USD) $Owner_{total}$ Owner pool total profit/loss (USD)
$ROwner_{toni}$ The ratio of $Owner_{total}$ and $Owner_{i}$ $ROwner_{roni}$ The ratio of $Owner_{r}$ and $Owner_{i}$
$ROwner_{roi}$ The ratio of ($Owner_{r}$--$Owner_{i}$) and $Owner_{i}$ $ROwner_{uonr}$ The ratio of $Owner_{u}$ and $Owner_{r}$
$ROwner_{uoni}$ The ratio of $Owner_{u}$ and $Owner_{i}$
$Impact_{min}$ Lowest pool impact of single owner’s action $RImpact_{minonavg}$ The ratio of $Impact_{min}$ and $Impact_{avg}$
$Impact_{max}$ Highest pool impact of single owner’s action $RImpact_{maxonavg}$ The ratio of $Impact_{max}$ and $Impact_{avg}$
$Impact_{avg}$ Average pool impact of all owners’ actions $RImpact_{minonmax}$ The ratio of $Impact_{min}$ and $Impact_{max}$
LPF $Age$ LP age ($=d$ if the pool still active) $IsAlive$ Whether the pool is still active
$Vol_{f}$ First day DEX volume $Pval_{f}$ LP’s total value after the first day
$Vol_{l}$ Last day DEX volume $Pval_{l}$ LP’s total value after the last day
$Vol_{max}$ Highest recorded DEX daily volume $Pval_{max}$ Highest recorded LP’s total value
$Vol_{min}$ Lowest recorded DEX daily volume $Pval_{min}$ Lowest recorded LP’s total value
$RVol_{fonl}$ The ratio of $Vol_{f}$ and $Vol_{l}$ $RPval_{fonl}$ The ratio of $Pval_{f}$ and $Pval_{l}$
$RVol_{fonmin}$ The ratio of $Vol_{f}$ and $Vol_{min}$ $RPval_{fonmin}$ The ratio of $Pval_{f}$ and $Pval_{min}$
$RVol_{fonmax}$ The ratio of $Vol_{f}$ and $Vol_{max}$ $RPval_{fonmax}$ The ratio of $Pval_{f}$ and $Pval_{max}$
$RVol_{lonmin}$ The ratio of $Vol_{last}$ and $Vol_{min}$ $RPval_{lonmin}$ The ratio of $Pval_{l}$ and $Pval_{min}$
$RVol_{lonmax}$ The ratio of $Vol_{last}$ and $Vol_{max}$ $RPval_{lonmax}$ The ratio of $Pval_{l}$ and $Pval_{max}$
$RVol_{minmax}$ The ratio of $Vol_{min}$ and $Vol_{max}$ $RPval_{minmax}$ The ratio of $Pval_{min}$ and $Pval_{max}$